### v2

# Pierce the Mists, Greet the Sky: Decipher Knowledge Overshadowing via Knowledge Circuit Analysis

Abstract
Large Language Models (LLMs), despite their remarkable capabilities, are hampered by hallucinations. A particularly challenging variant, knowledge overshadowing, occurs when one piece of activated knowledge inadvertently masks another relevant piece, leading to erroneous outputs even with high-quality training data. Current understanding of overshadowing is largely confined to inference-time observations, lacking deep insights into its origins and internal mechanisms during model training. Therefore, we introduce PhantomCircuit, a novel framework designed to comprehensively analyze and detect knowledge overshadowing. By innovatively employing knowledge circuit analysis, PhantomCircuit dissects the internal workings of attention heads, tracing how competing knowledge pathways contribute to the overshadowing phenomenon and its evolution throughout the training process. Extensive experiments demonstrate PhantomCircuit ’s effectiveness in identifying such instances, offering novel insights into this elusive hallucination and providing the research community with a new methodological lens for its potential mitigation.
[…]
tcb@breakable
Pierce the Mists, Greet the Sky:
Decipher Knowledge Overshadowing via Knowledge Circuit Analysis
Haoming Huang1,3, [^1] Co-first authors with equal contribution., Yibo Yan1,2, [^1] Co-first authors with equal contribution., Jiahao Huo1, Xin Zou1, Xinfeng Li4, Kun Wang4, Xuming Hu1,2, [^2] Corresponding author. 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology 3Shanghai Jiao Tong University   4Nanyang Technological University
1 Introduction
Large Language Models (LLMs) have witnessed explosive growth in recent years, demonstrating remarkable capabilities across a multitude of domains, including natural language understanding, generation, reasoning, and even cross-modal tasks Chang et al. (2024); Zhao et al. (2024); Yan et al. (2025a, 2024a); Xun et al. (2025). Their proficiency has catalyzed transformative advancements in various applications. However, a persistent challenge that tempers their widespread adoption and reliability is the phenomenon of hallucination. Broadly, hallucinations refer to instances where models generate content that is factually incorrect, nonsensical, or unfaithful to the provided source context, despite appearing coherent and fluent Rawte et al. (2023); Chakraborty et al. (2025); Wang et al. (2025).
Figure 1: Illustrative comparison of previous research with inference-time analysis (b) and observation-based explanation (c) vs our proposed PhantomCircuit (d) on knowledge overshadowing (a).
While substantial research has been dedicated to understanding the causes and detection of general hallucinations, a specific variant known as “knowledge overshadowing” warrants deeper investigation Zhang et al. (2024b, 2025a). This phenomenon is particularly perplexing because it can manifest even when models are trained on high-quality, meticulously curated pre-training corpora. Current understanding, primarily derived from inference-time observations, characterizes overshadowing as a scenario where, for a given query, one piece of activated knowledge inadvertently “overshadows” another relevant knowledge. This interference ultimately biases the model’s reasoning process, leading to a hallucinatory output, as illustrated in Figure 1 (a).
Nevertheless, existing explorations into knowledge overshadowing suffer from notable limitations. ❶ They predominantly focus on inference-time analysis, as shown in Figure 1 (b). While valuable for identifying the occurrence of overshadowing, such observations offer a surface-level understanding and often fall short of elucidating how these detrimental patterns are learned during the training phase. ❷ The explanations for overshadowing are often speculatively inferred from these observational outcomes rather than being rigorously investigated through dedicated interpretability tools that can probe the model’s internal decision-making mechanisms, as shown in Figure 1 (c). Consequently, a more comprehensive analytical framework is imperative to dissect this phenomenon from its origins to its manifestation.
To bridge this gap, we introduce PhantomCircuit, a novel framework designed to comprehensively analyze and detect the knowledge overshadowing phenomenon. Specifically, PhantomCircuit facilitates an in-depth examination of the evolution of overshadowing hallucinations throughout the training process, correlating their emergence and prevalence with core factors such as knowledge popularity, model size, and dataset size. Then, by leveraging knowledge circuit analysis as a key interpretability technique, we aim to trace the flow of information and the formation of knowledge representations within attention heads, thereby uncovering the internal mechanisms giving rise to overshadowing. Furthermore, we propose to optimize the number of edges within these circuits, thus alleviating the knowledge overshadowing. As illustrated in Figure 1 (d), our overall work aims to provide a clearer, mechanistic understanding and potential strategies for this elusive type of hallucination.
Our contributions can be summarized as follows:
  • We introduce PhantomCircuit, the first comprehensive framework designed to systematically analyze and detect knowledge overshadowing, delving into its mechanistic nature and evolution throughout model training.
  • We pioneer the use of knowledge circuit analysis to dissect the internal workings of attention heads, specifically elucidating how competing knowledge pathways contribute to the overshadowing phenomenon.
  • We conduct extensive experiments to demonstrate PhantomCircuit ’s efficacy in detecting knowledge overshadowing, offering novel insights and a new methodological lens for the research community.
2 Related Work
2.1 Hallucination Detection
Factuality hallucination detection, which aims to evaluate whether the output of LLMs aligns with real-world facts, typically involves either external fact-checking or internal uncertainty analysis Huang et al. (2025); Dang et al. (2024); Zheng et al. (2024); Zou et al. (2024); Zhou et al. (2024); Zhu et al. (2024). For instance, FACTSCORE Min et al. (2023) decomposes a generation into atomic facts and calculates the proportion that are supported by reliable knowledge sources. FACTOOL Chern et al. (2023), on the other hand, integrates multiple tools such as Google Search and Google Scholar to gather external evidence and assess the factuality of generated content. In contrast, methods like Chain-of-Verification Dhuliawala et al. (2023), probability-based assessments Kadavath et al. (2022); Zhang et al. (2024a), and uncertainty estimation approaches Varshney et al. (2023); Yao et al. (2023); Luo et al. (2023) rely on LLMs’ internal parametric knowledge or uncertainty signals to predict potential hallucinations. Among these efforts, knowledge overshadowing Zhang et al. (2025a) offers a novel perspective by modeling hallucination behavior from the perspective of knowledge representation, providing an efficient strategy for proactive prevention.
2.2 Knowledge Circuit Analysis
In the context of mechanistic interpretability Rai et al. (2024); Huo et al. (2024); Huang et al. (2024a), computations in Transformer-based language models are viewed as a connected directed acyclic graph encompassing components such as MLPs and attention layers Syed et al. (2023); Conmy et al. (2023); Huang et al. (2024b). A circuit refers to a sparse computational subgraph that significantly influences the model’s behavior on a specific task Olah et al. (2020); Elhage et al. (2021); Wang et al. (2022). Building on this, Yao et al. (2024) introduce the concept of knowledge circuits, hypothesizing that cooperation among components reveals implicit knowledge in LLMs. Further, Ou et al. (2025) explore how such circuits evolve during continual pre-training, providing insights into knowledge acquisition. To enable effective knowledge editing, CaKE Yao et al. (2024) proposes a Circuit-aware Knowledge Editing method that guides models to activate modified knowledge and form new reasoning pathways. In this paper, we analyze the phenomenon of knowledge overshadowing through the lens of knowledge circuits, contributing new perspectives to LLM hallucination detection.
See more related work in Appendix A.1.
3 Methodology
This section first defines the knowledge overshadowing phenomenon and its quantitative evaluation. Subsequently, we detail the PhantomCircuit framework, which encompasses methods for analyzing its training dynamics and for constructing and analyzing knowledge circuits to understand its internal mechanisms [^1] Our code will be available upon acceptance..
3.1 Knowledge Overshadowing
Knowledge overshadowing refers to a specific type of hallucination where less prevalent, subordinate knowledge is suppressed by high-frequency, dominant knowledge when both are associated with common background knowledge Zhang et al. (2025a).
Let $X_{dom}$ denote dominant knowledge entities and $X_{sub}$ denote subordinate knowledge entities, both potentially co-occurring with background knowledge $X_{bg}$. The core idea is that a strong learned pattern $X_{dom}\leftrightarrow X_{bg}$ can "over-generalize" where the model primarily associates $X_{dom}$ with $X_{bg}$.
Consequently, when the model encounters $X_{sub}$ with $X_{bg}$, denoted as $X_{sub}\leftrightarrow X_{bg}$), it may erroneously favor outputs related to $X_{dom}$ due to the stronger $X_{dom}\leftrightarrow X_{bg}$ pattern.
3.1.1 Knowledge Overshadowing Occurrence
When the input prompt is a knowledge pair composed of background knowledge and dominant knowledge, denoted as $P_{dom}=(X_{bg},X_{dom})$, and the model correctly generates the answer corresponding to the dominant knowledge, denoted as $Y_{dom}$, we consider this outcome, represented by the pair $(P_{dom},Y_{dom})$, as a successful recall of dominant knowledge.
When the input prompt is $P_{sub}$, but the model wrongly generates $Y_{dom}$, knowledge overshadowing occurs for this query-response instance, resulting in the $(P_{sub},Y_{dom})$.
3.1.2 Quantitative Indicators
Let $N_{dom}$ and $N_{sub}$ be the number of instances of $P_{dom}$ and $P_{sub}$ in the training dataset, respectively. The dataset $Z$ comprises a subset $Z_{dom}$ containing $N_{dom}$ instances of $P_{dom}$, and a subset $Z_{sub}$ containing $N_{sub}$ instances of $P_{sub}$.
During autoregressive generation tasks performed by the model, let $M_{sub}$ be the number of times when overshadowing instances $(P_{sub},Y_{dom})$ occur , and $M_{dom}$ be the number of times $(P_{dom},Y_{dom})$ occurs. Then, we can define the absolute extent of the knowledge overshadowing effect, the Absolute Overshadowing rate ($\mathcal{AO}$) and calculate it using $M_{sub}$ and $N_{sub}$,
$\mathcal{AO}=p(Y_{dom}|P_{sub})=\frac{M_{sub}}{N_{sub}}.$ (1)
To account for the model’s inherent performance and potential noise affecting the overshadowing rate, we also introduce $R_{dom}$ for dominant knowledge inputs:
$R_{dom}=p(Y_{dom}|P_{dom})=\frac{M_{dom}}{N_{dom}},$ (2)
which represents the recall rate for $P_{dom}$ query-response instances.
The Relative Overshadowing rate ($\mathcal{RO}$) is then defined as:
$\mathcal{RO}=\frac{\mathcal{AO}}{R_{dom}}=\frac{p(Y_{dom}|P_{sub})}{p(Y_{dom}|P_{dom})}.$ (3)
3.1.3 Overshadowing Influence Factors
Knowledge Popularity (P) is the fundamental cause of the knowledge overshadowing phenomenon and serves as its primary influencing factor. P is defined as the ratio of the capacities of $Z_{dom}$ to $Z_{sub}$, thus $P=N_{dom}/N_{sub}.$
Model Size (M), referring to the number of model parameters, also impacts knowledge overshadowing. A larger M generally implies stronger generalization capabilities, causing the model to rapidly generalize the $X_{dom}\leftrightarrow X_{bg}$ to instances involving $X_{sub}\leftrightarrow X_{bg}$ and exacerbating overshadowing.
In addition to the factors mentioned in Zhang et al. (2025a, 2024b), aiming to analyze the dynamic evolution of knowledge overshadowing during the training process, we extend our consideration to total number of tokens, the Dataset Size (D) in the training set. The average loss proportion of subordinate knowledge $\mathcal{LP}$ within an epoch is defined as $\mathcal{LP}=loss(P_{sub})/total\;loss,$ which also relates to the overshadowing.
3.2 Analysis Framework PhantomCircuit
Our proposed knowledge circuit-based overshadowing analysis framework involves the overshadowing dynamics analysis during training and circuit-based internal mechanism analysis.
3.2.1 Overshadowing Dynamic Analysis
Our framework provides a novel dynamic analysis of knowledge overshadowing during model training. By manipulating P, M, and D, we monitor $\mathcal{RO}$ across epochs. We focus on identifying the onset, duration, and recovery stages of overshadowing to understand their modulation by P, M, and D. Recognizing P as a key factor, we also explore how $\mathcal{LP}$ co-evolves with $\mathcal{RO}$ under these variables, aiming to uncover the role of $\mathcal{LP}$ in explaining overshadowing dynamics.
3.2.2 Knowledge Circuit Construction
We construct the knowledge circuit, a sparse computational subgraph $C\subseteq G$, where $G=(V,E)$ is the directed acyclic graph representation of LLMs, $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, $E$ represents the information flow between these components. Our goal is to identify a subgraph $C$ that is critical for recognizing the key component of given input prompt, particularly in knowledge overshadowing, is $\{X_{dom},X_{sub}\}$, the difference between $P_{dom}$ and $P_{sub}$. The adapted construction methods is similar to edge attribution patching Conmy et al. (2023), which involves:
Paired inputs. For a given $X_{bg}$, we create a pair of input prompts, the clean input $P_{sub}$ corresponding to our expected output $Y_{sub}$ in the overshadowing case, and the corrupt input $P_{dom}$ serving for the contrasting mentioned below.
Activation difference calculation. After running this pair of inputs through the model, we calculate the difference in activation values $\Delta A(v_{p}),\Delta A(v_{c})$ for the parent node $v_{p}$ and child node $v_{c}$ of each edge under these distinct inputs, $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corrupt}(v_{p}),$ where $A_{clean}(v_{p})$ and $A_{corrupt}(v_{p})$ are the activation of parent node when $P_{sub}$ as clean input and $P_{dom}$ as corrupt input.
Edge score calculation. The importance $S(e)$ of an edge $e=(v_{p},v_{c})$ is scored based on how patching $v_{p}$’s activation (using $\Delta A(v_{p})$) influences a metric $\mathcal{M}$, which assesses the model’s ability to correctly output $Y_{sub}$ rather than $Y_{dom}$ for $P_{sub}$ inputs. Using the Integrated Gradients, $S(e)$ is approximated as:
$S(e)\approx Exp\left[\Delta A(v_{p})\cdot\frac{\partial\mathcal{M}(Y_{sub}|P_{sub})}{\partial A(v_{p})}\right].$ (4)
Circuit construction. Edges with scores $|S(e)|$ below a threshold $\tau$ are pruned. The remaining subgraph forms the knowledge circuit $C$. See more details about construction in Appendix B.1
3.2.3 Circuit-based Analysis
PhantomCircuit mainly focuses on the attention heads in $C$ and follows these steps:
Node Attention Analysis. We identify high attention heads within $C$ by examining their attention scores and patterns, specifically their focus on $\{X_{dom},X_{sub}\}$.
Circuit Structure Analysis. We then trace the information flow of these high attention heads by identifying their parent and child nodes to understand their structural role. Nodes consistently retained in circuits built with different thresholds $\tau$ are also analyzed as key components.
Layer-wise Logit Evolution. Finally, using logit lens nostalgebraist (2020), we inspect the evolving output logits at layers associated with key nodes. This validates if their captured information contributes to the model’s prediction as expected.
3.2.4 Circuit-based Overshadowing Recovery
Inspired by the goal of knowledge circuits to maximize sensitivity to the distinguishing features between $X_{dom}$ and $X_{sub}$, we propose a circuit-based method to alleviate overshadowing. This involves optimizing the circuit structure by tuning the edge pruning threshold $\tau$ to obtain an optimal circuit, $C_{opt}$. The optimization is formulated as:
$\tau_{opt}=\operatorname*{arg\,max}_{\tau}\mathcal{M}(C_{opt}(\tau),P_{sub},Y_{sub}),$ (5)
where $\mathcal{M}$ measures the circuit’s ability to distinguish $\{X_{dom},X_{sub}\}$, This results in a $C_{opt}(\tau_{opt})$ which is expected to mitigate or eliminate the overshadowing effect for specific input prompts.
To automate the overshadowing recovery process and extend its applicability, we simplify the relative pointwise mutual information (R-PMI) method from Zhang et al. (2025a, 2024b) to identify $X_{sub}$ within $P_{sub}$. First, we obtain the top-$k$ next-token candidates, $V_{top}(P_{sub})$, by feeding $P_{sub}$ to the model. Then, we iteratively generate contrastive prompts $P^{\prime}_{sub}$ by masking (in our implementation, by deleting) each token $X^{\prime}_{sub}$ from $P_{sub}$. For each $P^{\prime}_{sub}$, we acquire its top-$k$ candidates $V_{top}(P^{\prime}_{sub})$. The R-PMI for each token $y_{i}$ in the intersection $V_{top}(P_{sub})\cap V_{top}(P^{\prime}_{sub})$ is calculated as:
$R\textnormal{-}PMI_{i}=\log\frac{p(y_{i}|P_{sub})}{p(y_{i}|P^{\prime}_{sub})}.$ (6)
Then, we sum the negative R-PMI values to get $S_{R\textnormal{-}PMI}=\sum\min(R\textnormal{-}PMI_{i},0)$. The $X^{\prime}_{sub}$ yielding the minimum $S_{R\textnormal{-}PMI}$ is identified as $X_{sub}$.
Furthermore, the $Y_{sub}$ is determined as the token $y_{i}$ from $V_{top}(P_{sub})$ whose average rank improves most significantly when non-subordinate components $X^{\prime}_{sub}$, often the $X_{bg}$, is masked. $Y_{dom}$ is identified as $y_{i}$ that has the highest average rank across all $P^{\prime}_{sub}$. For circuit construction, $X_{dom}$ within $P_{dom}$ is replaced by a generic placeholder like "something", as shown in Figure 5. Combining this streamlined approach enables broader application of our knowledge circuit-based overshadowing recovery. See more details in Appendix B.2
(a) Popularity (P)
(b) Model Size (M)
(c) Dataset Size (D)
(d) Popularity (P)
(e) Model Size (M)
(f) Dataset Size (D)
Figure 2: (a) $\sim$ (c) show the dynamic variation of $\mathcal{RO}$ relating to P, M and D in model training phase. Higher Knowledge Popularity (P) and Model Size (M) tend to result in an earlier onset, shorter duration, and quicker recovery from knowledge overshadowing. In contrast, a larger Dataset Size (D) also leads to an earlier onset but is associated with a slower recovery phase. (d) $\sim$ (f) show the duration of onset stage, high $\mathcal{RO}$ (> 90%) stage as well as recovery stage, and $\mathcal{RO}$’s rate of change during the onset and recovery stages.
4 Experiment
4.1 Experiment Setup
(a) M = 70M
(b) M = 410M
(c) M = 1.4B
Figure 3: The co-evolution between $\mathcal{RO}$ $\mathcal{LP}$ in different M. Early high overall loss and low $\mathcal{LP}$ leads to intensive optimization of $P_{dom}$ and $\mathcal{RO}$ rises up to near 100%. As training progresses, subordinate knowledge loss proportion ($\mathcal{LP}$) rises, shifting optimization focus to $P_{sub}$ errors, initiating $\mathcal{RO}$’s recovery phase, validated across models.
4.1.1 Dataset
Synthetic dataset. To investigate the dynamic characteristics and influencing factors of the $\mathcal{RO}$ during training under controlled conditions,and to minimize the complexities and semantic relationships inherent in natural language, we construct a synthetic dataset and train models from scratch.
Follow Zhang et al. (2025a), we first fix the length of $X_{bg}$ at 4 tokens and the lengths of $X_{dom},X_{sub},Y_{dom}$, $Y_{sub}$ at 1. For the dataset size (D), we experiment with 0.26 (D = 0.26M), 2.6 (D = 2.6M) and 26 million tokens (D = 26M). For popularity (P), we set P values as 5, 25, and 100.
Then, for a specific combination of P and D, there are several distinct groups to achieve the target D. Each group comprises P distinct $X_{dom}$ and a single $Y_{dom}$ for dominant knowledge, one $X_{sub}$ and $Y_{sub}$ for subordinate knowledge. The $X_{bg}$ is shared. Thus, there are P unique $\{P_{dom},Y_{dom}\}$ and one $\{P_{sub},Y_{sub}\}$ in a group. All the tokens are randomly sampled. See more details of our dataset in Appendix C.
Finetuning Dataset. We construct a finetuning dataset to evaluate circuit-base overshadowing recovery method. Utilizing virtual knowledge, we preserve the natural language semantics while avoiding prior knowledge editing that could interfere with the natural occurrence of overshadowing. For this dataset, we set P = 5 and D = 1M. The Appendix D shows the dynamics analysis based on finetuning dataset.
4.1.2 Models Evaluated
We employ models from the Pythia suite Biderman et al. (2023), specifically: Pythia-70M, Pythia-410M, Pythia-1.4B, and Pythia-2.8B, corresponding to model sizes (M) of M = 70M, 410M, 1.4B and 2.8B, respectively. Tokens randomly sampled to build synthetic dataset is from Pythia tokenizer.
4.1.3 Training
A uniform learning rate of $10^{-5}$ and batch size of 16 are used for both dataset. Training is conducted on NVIDIA A800 GPUs.
4.1.4 Evaluation
To measure $\mathcal{RO}$, we randomly sample 500 $P_{dom}$ and 500 $P_{sub}$ prompts for evaluation after each training epoch. The $\mathcal{LP}$ is recorded within each epoch. The results are shown in Figures 2 and 3.
Figure 4: The circuit-based analysis of proposed PhantomCircuit. (a) shows the average attention scores allocated to $\{X_{dom},X_{sub}\}$ across epoch. (b) focuses on the onset and recovery phases and shows that the appear and disappear of high attention heads (attention score greater than 0.2) attribute to the reduce and raise of $\mathcal{RO}$. (C) shows the logits and ranks of $Y_{sub},Y_{dom}$ across layers. The main structure of circuit (with 400 edges totally) and attention mode show that high attention head a5.h1 and MLP layer m5 contribute to the juncture of $Y_{sub}$ ’s rank. Solid lines denote direct information flow, while dashed lines indicate indirect flow in circuit structure map.
4.2 Main Result
Based on the experiments described above, using the circuit to analyze and optimize overshadowing, our investigation yields the following findings.
A higher value of P and M can lead to an earlier onset, shorter duration, and quicker recovery of the knowledge overshadowing. Distinctly, a larger D contributes to the earlier onset but also a slower recovery from overshadowing.
As shown in Figures 2(a) and 2(d) (M=70M, D=2.6M), increasing P (from 5 to 100) significantly shortens or even eliminates the onset phase. This is attributed to more prominent dominant patterns $X_{dom}\leftrightarrow X_{bg}$ being learned and generalized rapidly, even within the first epoch. The duration phase also decreases with higher P, because a larger P and a fixed D implies fewer groups of knowledge pairs, leading to less diversity of $P_{sub}$, allowing the model to learn all overshadowed $P_{sub}$ and recover from overshadowing more quickly.
Figures 2(b) and 2(e) (P=5, D=2.6M) demonstrate that larger models (M) exhibit shorter or absent onset phases, indicating an earlier occurrence of knowledge overshadowing. This is due to the stronger generalization capabilities of larger models, leading them to quickly learn and overgeneralize dominant patterns. However, larger models also show a shortened duration phase and a rapid decline in $\mathcal{RO}$ during recovery, suggesting enhanced capacity to differentiate $\{X_{dom},X_{sub}\}$, thus recovering faster despite earlier overshadowing.
In Figures 2(c) and 2(f) (P=5, M=70M) larger D leads to an earlier onset of overshadowing, with $\mathcal{RO}$ peaking sooner. This is because more data per epoch provides more iterations and exposure to the dominant pattern, accelerating its generalization. Conversely, the recovery phase is prolonged with larger datasets. The increased diversity of $P_{sub}$ in larger datasets requires more epochs for the model to learn all instances and recover.
Figure 5: Overshadowing recovery via optimized circuit. In the finetuned model, $X_{dom}\leftrightarrow X_{bg}$ causes the original full model to incorrectly predict $Y_{dom}$. First, we detect the $X_{sub},Y_{sub},Y_{dom}$ automatically by calculating the minimum $S_{R\textnormal{-}PMI}$ Then, optimizing the knowledge circuit by pruning edges to keep only key nodes enhances the attention on $\{X_{dom},X_{sub}\}$, enabling the recovery from overshadowing and facilitating correct $Y_{sub}$ prediction.
Notably, across various parameter combinations, $\mathcal{RO}$ often approaches 100% in the early training stages and finally recovers to 0%. We hypothesize this phenomenon stems from initial high-loss state, where optimization efforts disproportionately focus on reducing the larger loss contribution from $P_{dom}$. Our subsequent observations regarding $\mathcal{LP}$ corroborate this.
The dynamic nature of knowledge overshadowing arises from the co-evolution relationship between the $\mathcal{LP}$ and $\mathcal{RO}$. As depicted in Figure 3, in the early training stages, when the overall loss is high and the contribution from $P_{sub}$ is small, the optimization process tends to concentrate on the $P_{dom}$ and $\mathcal{RO}$ rapidly approaches 100%. However, as training progresses and $\mathcal{LP}$ begins to rise, eventually nearing its peak, $P_{sub}$ takes a substantial portion of the remaining loss. At this juncture, the model’s optimization efforts shift to focus on these errors from $P_{sub}$. This shift initiates the recovery phase, leading to a decline in $\mathcal{RO}$. Therefore, the insufficient optimization results in the overshadowing, which is consistent across varying M.
The occurrence of overshadowing in pretrained LLMs can be understood with dynamics analysis results above. Initially, large M and D promote the rapid generalization of $X_{dom}\leftrightarrow X_{bg}$ and overshadowing onset. The subsequent inadequate optimization of the $P_{sub}$ is exacerbated by the sheer scale and diversity of training data, which brings prolonged overshadowing effect and exceeded sharp recovery effect from large M (e.g., trillion tokens for training LlaMa-2-7B ).
The knowledge circuit’s attentional allocation to differences between dominant and overshadowed knowledge inputs dictates the extent of knowledge overshadowing. Figure 4 (a) shows the variation of circuit’s average attention scores on $\{X_{dom},X_{sub}\}$ throughout the training phase. The higher average attention alleviates overshadowing, while lower attention exacerbates it.
Figure 4 (b) shows the circuit dynamics across the onset and recovery phases of knowledge overshadowing. These bar plots show the attention scores of individual attention heads in a specific epoch, indicating that when the $\mathcal{RO}$ declines, some attention heads, defined as high attention heads, exhibiting high focus on $\{X_{dom},X_{sub}\}$ emerge within the circuit. The threshold for high attention is set at 0.2 for the length of $X_{sub}$ is one fifth of the length of input prompt in synthetic dataset. Conversely, when knowledge overshadowing intensifies, a subset of these critical attention heads tends to disappear from the circuit.
Furthermore, by focusing on the circuit’s internal mechanisms and structure within a specific epoch, as shown in Figure 4 (c), we leverage the circuit-based analysis of our proposed PhantomCircuit. First, according to the logits and ranks of $Y_{dom}$ and $Y_{sub}$ across layers, the 5th layer is a juncture where the rank of $Y_{sub}$ exceeds $Y_{dom}$. Subsequently, by focusing on the circuit structure, we identify the internal mechanisms driving this juncture. A high attention head, a5.h1 (layer 5), crucially channels information to the subsequent MLP layer, m5. The a5.h1 also appears to inherit its high attention state from earlier layers, mediated by the layer 4 MLP (m4), leading it to not only continue elevating the logits for $Y_{sub}$ but also to attenuate the previously rapid growth of $Y_{dom}$’s logits, thereby facilitating the observed rank reversal. The attention map on the right confirms its high focus on $\{X_{dom},X_{sub}\}$.
Knowledge circuit-guided optimization represents a promising strategy to mitigate the knowledge overshadowing effect. Figure 5 illustrates our findings. We first finetune the model on finetuning dataset. During the recovery phase, we randomly choose some $P_{sub}$. Feeding $P_{sub}$ to the original full model (M = 410M), the prediction is still $Y_{dom}$ instead of expected $Y_{sub}$, and shows a significant gap in probability as well.
We then detect the $Y_{sub},Y_{dom}$ and $X_{sub}$ for chosen $P_{sub}$ automatically, replace the $X_{sub}$ of placeholder token "something". With these components, we obtain the $P_{dom}$ as corrupt input and $P_{sub}$ as clean input to construct a knowledge circuit. A golden section search algorithm is employed to determine the optimal number of edges for building $C_{opt}$. The optimized circuit structure map shows that some attention heads are pruned, which are often low attention heads, or exhibit no significant attentional pattern towards the core task-relevant information. Retained high attention heads are key to differentiating $P_{dom}$ from $P_{sub}$ which give significant attention to $\{X_{dom},X_{sub}\}$. Some low attention heads also remain, implying that even in the circuit, processing background knowledge $X_{bg}$ and linking it to the distinctive elements $X_{sub}$ is crucial for correct inference. The performance of the optimized circuit $C_{opt}$ is then evaluated by feeding it the clean input $P_{sub}$, while $P_{dom}$ serves as the baseline for contrast. Finally, the circuit successfully produces $Y_{sub}$, demonstrating the elimination of the overshadowing effect.
Future work will enhance the circuit optimization metric $\mathcal{M}$ by incorporating $Y_{sub}$’s absolute logit alongside the logit difference with $Y_{dom}$ for more effective guidance. Developing a comprehensive evaluation framework for circuit-based recovery is also crucial. These steps will evolve PhantomCircuit into an integrated platform for efficient analysis and robust optimization of knowledge overshadowing.
5 Conclusion
This paper investigates hallucinations in LLMs caused by knowledge overshadowing, and introduces PhantomCircuit, a novel knowledge circuit-based analysis framework. PhantomCircuit first analyzes the training dynamics of overshadowing, finding that dominant knowledge popularity, model size, and dataset size critically shape the onset, duration, and recovery of overshadowing. Apart from that, the persistent overshadowing in pretrained models stems from inadequately optimized subordinate knowledge loss. By analyzing knowledge circuits, we find that changes in critical attention heads’ focus on subordinate knowledge directly correlate with the recovery or onset of overshadowing. Finally, optimizing these knowledge circuits presents a promising strategy for mitigating knowledge overshadowing.
Limitations
Despite the insights provided by PhantomCircuit, this study has several limitations that open avenues for future research:
  1. 1.
    The dynamic analysis of knowledge circuits throughout training is computationally intensive, potentially hindering scalability to very large models or extensive training. We aim to develop more computationally efficient techniques for approximating circuit evolution, such as checkpoint-based analysis or lightweight probing.
  2. 2.
    This study concentrates on a specific type of knowledge overshadowing, leaving more complex or subtle interference patterns unaddressed. Future work will broaden PhantomCircuit’s scope to investigate a wider range of overshadowing phenomena, including those in multi-hop reasoning.
  3. 3.
    Future efforts will focus on evolving our instance-specific circuit optimizations into a generalized mitigation toolkit, supported by a comprehensive evaluation framework. Key improvements will target the precision of automated overshadowed knowledge identification and the broader efficacy of circuit-based interventions. Ultimately, we aim to develop PhantomCircuit as a robust platform for both in-depth analysis and effective, generalizable overshadowing mitigation.
Appendix A More Related Work
A.1 Large Language Models
First proposed by Brown et al. (2020), Transformer-based auto-regressive LLMs have demonstrated strong performance across a variety of NLP tasks, including question answering Yue (2025), in-context learning Dong et al. (2022), and analogical reasoning Webb et al. (2023). Pretrained on large-scale text corpora, LLMs have acquired extensive real-world knowledge from web sources. As a result, models such as InternLM2.5 Cai et al. (2024), Qwen2.5 Team (2024), and LLaMA3.3 Grattafiori et al. (2024) have shown excellent performance on world knowledge benchmarks Suzgun et al. (2022). Therefore, over the past year, LLMs have demonstrated remarkable capabilities in understanding-related tasks across various fields Kim et al. (2024); Nam et al. (2024); Yan et al. (2024c); Yan and Lee (2024); Yan et al. (2024b); Dong et al. (2025); Su et al. (2025).
Recently, there has been a growing trend toward enhancing LLMs’ reasoning capabilities on complex tasks Guo et al. (2025); Jaech et al. (2024) by generating long Chain-of-Thoughts (CoTs), with reinforcement learning (RL) emerging as an effective tool to encourage this behavior Li et al. (2025); Trung et al. (2024). Recently, there have also been efforts to explore collaboration between LLMs to enhance their reasoning abilities Zhang et al. (2025b); Putta et al. (2024); Masterman et al. (2024); Yan et al. (2025b); Chu et al. (2025).
Despite these advancements, existing LLMs still suffer from factual hallucinations in practice Pan et al. (2025); Asgari et al. (2025), with knowledge overshadowing identified as a primary contributing factor Zhang et al. (2024b). While existing interoperability works make great efforts on the mechanism of LLM training and generating Zhao et al. (2024), most of them solely focus on isolated model versions like GPT2 Wang et al. (2023) and LLaMA2 Wendler et al. (2024); Tang et al. (2024).
In this paper, we utilize the Pythia suite Biderman et al. (2023) to investigate the evolution and underlying mechanisms of knowledge overshadowing across models of varying sizes: 70M, 410M, 1.4B, and 2.8B parameters. Sharing a unified architecture, this model suite eliminates design variability, thereby providing clearer and more reliable insights into the scaling behavior of the knowledge overshadowing phenomenon in LLMs.
Appendix B PhantomCircuit Details
B.1 Circuit Construction
Knowledge circuit is as a sparse computational subgraph within the LLMs. The construction of such a circuit involves identifying and retaining the most influential components (nodes, including MLPs and attention heads) and connections (edges) while pruning less critical ones.
We adapted the optimized circuit construction method provided by Yao et al. (2024). The process begins by representing the LLM as a directed acyclic graph (DAG), $G=(V,E)$, where $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, and $E$ represents the information flow between these components. The goal is to identify a subgraph $C\subseteq G$ that is critical for recognize the key component of a given input prompt, particularly in knowledge overshadowing, is $X_{dom}$ and $X_{sub}$, the difference between $P_{dom}$ and $P_{sub}$.
The adapted construction method is similar to edge attribution patching(EAP) Conmy et al. (2023), which involves:
  1. 1.
    Paired Inputs: For a given background $X_{bg}$, we create two primary input prompts: $P_{dom}=(X_{bg},X_{dom})$ and $P_{sub}=(X_{bg},X_{sub})$. We also consider a "corrupted" version of $P_{sub}$, which could be $P_{dom}$ itself or another prompt designed to elicit $Y_{dom}$. Let’s denote the "clean" input as $P_{clean}$ (typically $P_{sub}$) and the "corrupted" input as $P_{corr}$ (designed to lead to $Y_{dom}$).
  2. 2.
    Activation Difference Calculation: We run both $P_{clean}$ and $P_{corr}$ through the model. For each node $v\in V$ that is a potential parent in an edge, we record its output activation. The difference in activations between the clean and corrupted runs for a node $v_{p}$ (parent) is denoted as $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corr}(v_{p})$.
  3. 3.
    Edge Scoring via Gradient-based Attribution: To score an edge $e=(v_{p},v_{c})$ (from parent $v_{p}$ to child $v_{c}$), we focus on how patching the activation from $v_{p}$ (i.e., using $A_{clean}(v_{p})$ instead of $A_{corr}(v_{p})$ when $P_{corr}$ is the main input) affects a chosen metric $\mathcal{M}$. This metric $\mathcal{M}$ is designed to measure the model’s tendency towards generating $Y_{sub}$ versus $Y_{dom}$ when the input is $P_{sub}$. A common choice for $\mathcal{M}$ could be the logit difference between $Y_{sub}$ and $Y_{dom}$ at the final layer, or a metric related to our Relative Overshadowing rate ($RO$).
    The score $S(e)$ for an edge $e$ can be approximated by the product of the activation difference from its parent node and the gradient of the metric $\mathcal{M}$ with respect to the input of its child node, when the child node receives the "clean" activation from the parent while other inputs are "corrupted":
    $S(e)\approx\mathbb{E}_{P_{sub}}\left[\Delta A(v_{p})\cdot\frac{\partial\mathcal{M}(Y_{target}|P_{sub})}{\partial A_{input}(v_{c})}\right]$
    where $Y_{target}$ is ideally $Y_{sub}$. The expectation $\mathbb{E}$ is taken over instances of $P_{sub}$ in our evaluation set $Z_{sub}$. In practice, methods like Integrated Gradients (IG) are often used to refine this attribution by integrating gradients along a path from a baseline (corrupted) input to the actual (clean) input.
  4. 4.
    Circuit Pruning: Based on the calculated scores $S(e)$, edges with scores below a certain threshold $\tau$, or alternatively, edges outside the top-N highest scores, are pruned from the graph $G$. The remaining nodes and edges form the knowledge circuit $C_{sub}$.
    $C_{sub}=(V_{sub},E_{sub})$
    where $E_{sub}=\{e\in E\mid|S(e)|\geq\tau\}$ (or top-N criterion) and $V_{sub}$ consists of nodes connected by edges in $E_{sub}$.
    This constructed circuit $C_{sub}$ is then analyzed to understand how dominant knowledge $K_{dom}$ might overshadow $K_{sub}$ by examining the attentional features and information flow within it, especially when processing $P_{sub}$.
B.2 Automated Component Identification for Recovery
Identifying the Overshadowed Component $X_{sub}^{*}$. A critical precursor to effective circuit-based recovery is the precise identification of the specific component $X_{sub}^{*}$ within the subordinate prompt $P_{sub}$ that is being overshadowed. This is achieved by adapting the Relative Pointwise Mutual Information (R-PMI) based methodology from Zhang et al. (2025a, 2024b). The process involves:
Iteratively generating contrastive prompts $P^{\prime}_{sub}$ by deleting each candidate token $X^{\prime}_{sub}$ (a potential overshadowed component) from the original $P_{sub}$.
For each pair $(P_{sub},P^{\prime}_{sub})$, calculating the R-PMI for tokens $y_{i}$ in the intersection of their top-$k$ next-token candidate sets, $V_{top}(P_{sub})\cap V_{top}(P^{\prime}_{sub})$, using
$R\textnormal{-}PMI(y_{i};P_{sub},P^{\prime}_{sub})={}$ $\log P(y_{i}|P_{sub})$
$-\log P(y_{i}|P^{\prime}_{sub}).$
Summing only the negative R-PMI values to obtain
$S_{R\textnormal{-}PMI^{-}}(P_{sub},P^{\prime}_{sub})=\sum\min(R\textnormal{-}PMI(y_{i}),0).$
The $X^{\prime}_{sub}$ that yields the minimum (most negative) $S_{R\textnormal{-}PMI^{-}}$ is identified as the primary overshadowed component, $X_{sub}^{*}$. This selection is based on the rationale that removing the true $X_{sub}^{*}$ most strongly exposes the model’s bias towards outputs favored by the dominant knowledge pattern.
Identifying Target Subordinate Output $Y_{sub}$. The intended subordinate output $Y_{sub}$ is identified by assessing which token from $V_{top}(P_{sub})$ (the top-k candidates for the original prompt $P_{sub}=(X_{bg},X_{sub})$) exhibits the most significant improvement when the overshadowing influence of background knowledge ($X_{bg}$) or other non-subordinate components is mitigated. Specifically, we generate contrastive prompts $P^{\prime}_{sub}$ by masking or altering components of $X_{bg}$ (or other identified non-subordinate elements that contribute to the $X_{bg}\leftrightarrow Y_{dom}$ association) within $P_{sub}$. $Y_{sub}$ is then the token $y_{i}\in V_{top}(P_{sub})$ that shows the most substantial rank improvement (or largest increase in log probability) in these modified prompts $P^{\prime}_{sub}$ compared to its rank in the original $P_{sub}$. This rank elevation signifies the “unmasking” of the true subordinate answer as the dominant, overshadowing associations are weakened.
Identifying Dominant Output $Y_{dom}$. The dominant output $Y_{dom}$ is identified as the token that maintains the highest average rank across all contrastive prompts $P^{\prime}_{sub}$ generated by deleting different candidate tokens $X^{\prime}_{sub}$ from $P_{sub}$. This token represents the model’s most consistent, default output tendency when specific subordinate cues are variably weakened, likely reflecting the pervasive influence of dominant knowledge associated with the background $X_{bg}$.
[…]
Some more circuit-based overshadowing recovery cases are shown in Table 1.
Table 1: Circuit-based overshadowing recovery cases
{tblr}
[…]
Rank 1: Logit: 15.42 Prob: 14.98% Token: |Time| Rank 1: Logit: 19.75 Prob: 20.42% Token: |Tem|
Rank 2: Logit: 14.21 Prob: 4.47% Token: |Sp| Rank 2: Logit: 17.34 Prob: 1.84% Token: |Filter|
Rank 3: Logit: 14.05 Prob: 3.83% Token: |T| Rank 3: Logit: 16.40 Prob: 0.72% Token: |E|
Rank 4: Logit: 13.84 Prob: 3.10% Token: |Custom| Rank 4: Logit: 16.29 Prob: 0.64% Token: |Sp|
[…]
Rank 1: Logit: 16.89 Prob: 17.65% Token: |E| Rank 1: Logit: 16.47 Prob: 10.39% Token: |Ps|
Rank 2: Logit: 15.35 Prob: 3.76% Token: |Ps| Rank 2: Logit: 15.47 Prob: 3.84% Token: |Wave|
Rank 3: Logit: 15.29 Prob: 3.57% Token: |St| Rank 3: Logit: 15.31 Prob: 3.28% Token: |Energy|
Rank 4: Logit: 15.15 Prob: 3.09% Token: |emitter| Rank 4: Logit: 14.89 Prob: 2.15% Token: |emitter|
[…]
Rank 1: Logit: 48.68 Prob: 33.09% Token: |St| Rank 1: Logit: 44.01 Prob: 27.96% Token: |Transfer|
Rank 2: Logit: 46.66 Prob: 4.38% Token: |Ada| Rank 2: Logit: 42.38 Prob: 5.48% Token: |St|
Rank 3: Logit: 46.37 Prob: 3.28% Token: |Flow| Rank 3: Logit: 42.24 Prob: 4.75% Token: |Energy|
Rank 4: Logit: 46.06 Prob: 2.41% Token: |Emit| Rank 4: Logit: 41.64 Prob: 2.61% Token: |Mi|
[…]
Rank 1: Logit: 38.34 Prob: 6.67% Token: |St| Rank 1: Logit: 34.24 Prob: 35.19% Token: |Field|
Rank 2: Logit: 35.87 Prob: 0.56% Token: |Fl| Rank 2: Logit: 31.41 Prob: 2.08% Token: |Output|
Rank 3: Logit: 35.82 Prob: 0.54% Token: |Trans| Rank 3: Logit: 30.97 Prob: 1.34% Token: |Har|
Rank 4: Logit: 33.51 Prob: 0.05% Token: |Un| Rank 4: Logit: 30.84 Prob: 1.18% Token: |Fl|
Appendix C Dataset Details
C.1 Detailed Synthetic Dataset Construction
The synthetic dataset was constructed through the following steps to ensure controlled conditions for analyzing knowledge overshadowing dynamics:
Fixing text lengths. For all generated data instances, consistent token lengths are maintained. The background knowledge ($X_{bg}$) was set to a length of 4 tokens. All other core components, namely the dominant knowledge entity ($X_{dom}$), subordinate knowledge entity ($X_{sub}$), dominant output ($Y_{dom}$), and subordinate output ($Y_{sub}$), are each set to a length of 1 token.
Dataset generation for specific D and P Combinations. For each defined combination of total dataset size (D) and knowledge popularity (P), the dataset was built as follows: The dataset comprises multiple distinct groups of knowledge instances. Each group consists of P+1 knowledge prompts: a set of P dominant knowledge prompts $\{P_{dom}^{1},P_{dom}^{2},\dots,P_{dom}^{P}\}$ and one subordinate knowledge prompt $P_{sub}$. Within each group:
  • For the P dominant prompts, the actual dominant knowledge entities $\{X_{dom}^{1},X_{dom}^{2},\dots,X_{dom}^{P}\}$ are all unique. However, they all share the same background knowledge component ($X_{bg}$) and are associated with the same dominant output ($Y_{dom\_g}$). Thus, each $P_{dom}^{i}=(X_{bg},X_{dom}^{i})$ is paired with $Y_{dom}^{g}$.
  • The single subordinate prompt $P_{sub}=(X_{bg},X_{sub})$ uses the same background knowledge $X_{bg}$ as the dominant prompts in that group. However, its subordinate knowledge entity $X_{sub}$ is distinct from all $X_{dom}^{i}$ entities in that group, and its corresponding output $Y_{sub}$ is distinct from any $Y_{dom}$ in group.
This structure creates a group: $\{(P_{dom}^{1},Y_{dom}^{1}),\dots,((P_{dom}^{P},Y_{dom}^{P}),(P_{sub},Y_{sub})\}$.
Multiple such groups are generated. All tokens for $X_{bg},X_{dom}^{i},X_{sub},Y_{dom},Y_{sub}$ within each group, and across different groups, are randomly sampled from the Pythia tokenizer vocabulary, ensuring no overlap between the core entities of different groups. This process was repeated until the total number of tokens in the dataset reached the target size D.
Cases illustration. We illustrate some groups for P=5 dataset in Table 2. We directly show token id.
Table 2: Illustrative examples from the synthetic dataset (P=5). Each data entry is a row, with fine lines separating entries within a group. Token IDs are shown.
{tblr}
[…]
$[10030,16936,1050,10565]$ $24327$ $20730$
$[10030,16936,1050,10565]$ $4619$ $20730$
$[10030,16936,1050,10565]$ $5137$ $20730$
$[10030,16936,1050,10565]$ $785$ $20730$
$[10030,16936,1050,10565]$ $18941$ $3519$
[…]
$[17026,8837,3802,28741]$ $3530$ $1077$
$[17026,8837,3802,28741]$ $11948$ $1077$
$[17026,8837,3802,28741]$ $2028$ $1077$
$[17026,8837,3802,28741]$ $9389$ $1077$
$[17026,8837,3802,28741]$ $25814$ $5374$
[…]
$[18131,14501,21161,311]$ $778$ $7790$
$[18131,14501,21161,311]$ $18762$ $7790$
$[18131,14501,21161,311]$ $28591$ $7790$
$[18131,14501,21161,311]$ $28981$ $7790$
$[18131,14501,21161,311]$ $8447$ $6129$
C.2 Finetuning dataset
For the finetuning dataset, we utilized the Qwen-Long API to generate instances of virtual knowledge. This generated data subsequently underwent manual review to identify and remove any instances that are overly repetitive or semantically too similar, ensuring a degree of diversity, resulted in D = 1M.
A key distinction from the synthetic dataset construction is that we did not strictly control token lengths for each component in this dataset. Instead of randomly sampled token IDs, the finetuning dataset consists of actual linguistic statements that, while syntactically and semantically coherent, represent virtual (i.e., fabricated but plausible) knowledge. The underlying pattern of dominant and subordinate knowledge construction, however, mirrors that of the synthetic dataset.
As an example of this dataset, we set the knowledge popularity P=5. Some illustrative cases from the dataset are shown in Table 3.
Table 3: Illustrative examples from the Finetuning Dataset (P=5). Each data entry is a row, with fine lines separating entries within a group.
{tblr}
[…]
Analysis of the Starfire Crystal Engine reveals primary energy output idle cycle Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output phase synchronicity Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output null gravity Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output core temperature Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output cold ignition Quantum Flux
[…]
Observation of the Chronos Orb’s temporal distortion signature indicates near Leyline Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates while attuned Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates during solstice Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates etherium interaction Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates sudden impact Temporal Anomaly
[…]
Study of the Xylosian biology reveals primary nutrient absorption method after sonic bath Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method during digestion Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method high pressure Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method thermal vent Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method xenoflora consumption Crystalline Excretion
(a) The $\mathcal{RO}$ during training phase of the different finetuning dataset size (D).
(b) The co-evolution of $\mathcal{LP}$ and $\mathcal{RO}$ during training phase on finetuning dataset.
Figure 6: The dynamics analysis of knowledge overshadowing in finetuning dataset.
Appendix D More Dynamics Analysis on Finetuning Dataset
In addition to validating the efficacy of our circuit-based overshadowing recovery method, the finetuning dataset serves a dual purpose. We also leverage it to empirically verify our conclusions regarding the training dynamics of knowledge overshadowing, specifically concerning the impact of Dataset Size (D). Consistent with our dynamic analysis findings, we investigate whether a larger D indeed correlates with a slower recovery rate from knowledge overshadowing and a prolonged duration of the hallucination effect. To this end, we conduct experiments on the finetuning dataset by fixing Knowledge Popularity at P=5 and Model Size at M=70M, while varying D across values of 0.1M, 0.5M, and 1M tokens. The results, as depicted in Figure 6(a), corroborate this relationship. Furthermore, under the specific configuration of P=5, M=70M, and D=0.5M on the finetuning dataset, we re-examine the interplay between the loss proportion of subordinate knowledge ($\mathcal{LP}$) and the relative overshadowing rate ($\mathcal{RO}$). As shown in Figure 6(b), the observations again support the hypothesis that insufficient optimization of subordinate knowledge contributes to the persistence of knowledge overshadowing.
It is noteworthy that distinct behaviors are observed when comparing the finetuning dataset to the synthetic dataset. Firstly, the recovery from overshadowing on the finetuning dataset is generally slower than on the synthetic dataset for same D. This can be attributed to the richer semantic relationships and greater complexity inherent in the natural language of the finetuning data, which presents a more challenging learning task.
Secondly, we observe that the finetuning dataset exhibits a minimal or absent onset phase for knowledge overshadowing, where $\mathcal{RO}$ typically rise. This is because finetuning commences from a pretrained model, which has already moved beyond the initial epochs of chaotic, random predictions. Consequently, the model can very rapidly generalize strong association patterns present in the finetuning data. Moreover, the diverse and varied forms of data within the finetuning set may act akin to a beneficial noise signal, prompting the model to pay closer attention to distinguishing features and differences. This inherent data diversity can help preemptively mitigate or even eliminate the early onset stage of knowledge overshadowing that might otherwise be observed.

### v4

@@
 Therefore, we introduce PhantomCircuit, a novel framework designed to comprehensively analyze and detect knowledge overshadowing.
-By innovatively employing knowledge circuit analysis, PhantomCircuit dissects the internal workings of attention heads, tracing how competing knowledge pathways contribute to the overshadowing phenomenon and its evolution throughout the training process.
+By innovatively employing knowledge circuit analysis, PhantomCircuit dissects the function of key components in the circuit and how the attention pattern dynamics contribute to the overshadowing phenomenon and its evolution throughout the training process.
 Extensive experiments demonstrate PhantomCircuit ’s effectiveness in identifying such instances, offering novel insights into this elusive hallucination and providing the research community with a new methodological lens for its potential mitigation.
+Our code can be found in https://github.com/halfmorepiece/PhantomCircuit.
 […]
@@
 Haoming Huang1,3, [^1] Co-first authors with equal contribution., Yibo Yan1,2, [^1] Co-first authors with equal contribution., Jiahao Huo1, Xin Zou1, Xinfeng Li4, Kun Wang4, Xuming Hu1,2, [^2] Corresponding author.
-1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology 3Shanghai Jiao Tong University   4Nanyang Technological University
+1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology 3Shanghai Jiao Tong University   4Nanyang Technological University {hmhuang04, yanyibo70}@gmail.com, {xuminghu}@hkust-gz.edu.cn
 1 Introduction
@@
 Further, Ou et al.
+(2025); Zucchet et al.
+(2025); Hakimi et al.
 (2025) explore how such circuits evolve during continual pre-training, providing insights into knowledge acquisition.
@@
 See more related work in Appendix A.1.
+Figure 2: A group of imbalanced natural language data of knowledge overshadowing.
+The $X_{bg}$, $X_{dom}$ and $X_{sub}$ are background, dominant and subordinate knowledge respectively.
+The $P_{dom}$ and $P_{sub}$ are prompt with dominant and subordinate knowledge respectively, their corresponding answers are $Y_{dom}$ and $Y_{sub}$.
 3 Methodology
@@
 (2025a).
-Let $X_{dom}$ denote dominant knowledge entities and $X_{sub}$ denote subordinate knowledge entities, both potentially co-occurring with background knowledge $X_{bg}$.
+As shown in Figure 2, let $X_{dom}$ denote dominant knowledge entities and $X_{sub}$ denote subordinate knowledge entities, both potentially co-occurring with background knowledge $X_{bg}$.
 The core idea is that a strong learned pattern $X_{dom}\leftrightarrow X_{bg}$ can "over-generalize" where the model primarily associates $X_{dom}$ with $X_{bg}$.
-Consequently, when the model encounters $X_{sub}$ with $X_{bg}$, denoted as $X_{sub}\leftrightarrow X_{bg}$), it may erroneously favor outputs related to $X_{dom}$ due to the stronger $X_{dom}\leftrightarrow X_{bg}$ pattern.
+Consequently, when the model encounters $X_{sub}$ with $X_{bg}$, denoted as $X_{sub}\leftrightarrow X_{bg}$, it may erroneously favor outputs related to $X_{dom}$ due to the stronger $X_{dom}\leftrightarrow X_{bg}$ pattern.
 3.1.1 Knowledge Overshadowing Occurrence
@@
 When the input prompt is $P_{sub}$, but the model wrongly generates $Y_{dom}$, knowledge overshadowing occurs for this query-response instance, resulting in the $(P_{sub},Y_{dom})$.
+In the case shown in Figure 2, if overshadowing occurs, the model will incorrectly answer "Beijing" to the input prompt "Where is the most important economic center of China?".
 3.1.2 Quantitative Indicators
-Let $N_{dom}$ and $N_{sub}$ be the number of instances of $P_{dom}$ and $P_{sub}$ in the training dataset, respectively.
-The dataset $Z$ comprises a subset $Z_{dom}$ containing $N_{dom}$ instances of $P_{dom}$, and a subset $Z_{sub}$ containing $N_{sub}$ instances of $P_{sub}$.
+Let $N_{dom}$ be the number of instances of the dominant knowledge prompt $P_{dom}$ and $N_{sub}$ be the number of instances of the subordinate knowledge prompt $P_{sub}$ in a group of data, respectively.
+In the case shown in Figure 2, $N_{dom}$ and $N_{sub}$ are 2 and 1, respectively.
+The entire dataset might contain multiple different groups of imbalanced data.
 During autoregressive generation tasks performed by the model, let $M_{sub}$ be the number of times when overshadowing instances $(P_{sub},Y_{dom})$ occur, and $M_{dom}$ be the number of times $(P_{dom},Y_{dom})$ occurs.
@@
 Knowledge Popularity (P) is the fundamental cause of the knowledge overshadowing phenomenon and serves as its primary influencing factor.
-P is defined as the ratio of the capacities of $Z_{dom}$ to $Z_{sub}$, thus $P=N_{dom}/N_{sub}.$
+P is defined as the ratio of $N_{dom}$ to $N_{sub}$ in a group,
+$P=\frac{N_{dom}}{N_{sub}}.$ (4)
+In the case of Figure 2, $P=2$. A higher value of P signifies a more severe data imbalance.
 Model Size (M), referring to the number of model parameters, also impacts knowledge overshadowing.
@@
 (2025a, 2024b), aiming to analyze the dynamic evolution of knowledge overshadowing during the training process, we extend our consideration to total number of tokens, the Dataset Size (D) in the training set.
-The average loss proportion of subordinate knowledge $\mathcal{LP}$ within an epoch is defined as $\mathcal{LP}=loss(P_{sub})/total\;loss,$ which also relates to the overshadowing.
+Furthermore, the average loss proportion of subordinate knowledge $\mathcal{LP}$ within an epoch during training is defined as
+$\mathcal{LP}=\frac{loss(P_{sub})}{total\;loss},$ (5)
+which measures the allocation of the model’s efforts for the optimization of subordinate knowledge and recovery from knowledge overshadowing.
 3.2 Analysis Framework PhantomCircuit
@@
 3.2.2 Knowledge Circuit Construction
-We construct the knowledge circuit, a sparse computational subgraph $C\subseteq G$, where $G=(V,E)$ is the directed acyclic graph representation of LLMs, $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, $E$ represents the information flow between these components.
+We construct the knowledge circuit, a sparse computational subgraph $C\subseteq G$, where $G=(V,E)$ is the directed acyclic graph representation of LLMs.
+The node set $V$ encompasses input embeddings, attention heads, MLP layers, and output logits.
+The edge set $E$ represents the information flow between these components.
 Our goal is to identify a subgraph $C$ that is critical for recognizing the key component of given input prompt, particularly in knowledge overshadowing, is $\{X_{dom},X_{sub}\}$, the difference between $P_{dom}$ and $P_{sub}$.
-The adapted construction methods is similar to edge attribution patching Conmy et al.
+The adapted construction method is similar to Edge Attribution Patching with Integrated Gradients (EAP-IG) Hanna et al.
 (2024), which involves:
 Paired inputs.
-For a given $X_{bg}$, we create a pair of input prompts, the clean input $P_{sub}$ corresponding to our expected output $Y_{sub}$ in the overshadowing case, and the corrupt input $P_{dom}$ serving for the contrasting mentioned below.
+For a given $X_{bg}$, we create a pair of input prompts, the clean input $P_{sub}$ corresponding to our expected output $Y_{sub}$ in the overshadowing case, and the corrupt input $P_{dom}$ serving for the contrast mentioned below.
 Activation difference calculation.
-After running this pair of inputs through the model, we calculate the difference in activation values $\Delta A(v_{p}),\Delta A(v_{c})$ for the parent node $v_{p}$ and child node $v_{c}$ of each edge under these distinct inputs, $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corrupt}(v_{p}),$ where $A_{clean}(v_{p})$ and $A_{corrupt}(v_{p})$ are the activation of parent node when $P_{sub}$ as clean input and $P_{dom}$ as corrupt input.
+After running this pair of inputs through the model, we calculate the difference in activation values $\Delta A(v_{p}),\Delta A(v_{c})$ for the parent node $v_{p}$ and child node $v_{c}$ of each edge under these distinct inputs, $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corrupt}(v_{p}),$ where $A_{clean}(v_{p})$ and $A_{corrupt}(v_{p})$ are the activations of the parent node for the clean input $P_{sub}$ and the corrupt input $P_{dom}$, respectively.
 Edge score calculation.
 The importance $S(e)$ of an edge $e=(v_{p},v_{c})$ is scored based on how patching $v_{p}$’s activation (using $\Delta A(v_{p})$) influences a metric $\mathcal{M}$, which assesses the model’s ability to correctly output $Y_{sub}$ rather than $Y_{dom}$ for $P_{sub}$ inputs.
-Using the Integrated Gradients, $S(e)$ is approximated as:
+Using the Integrated Gradients (IG), $S(e)$ is approximated as:
 $S(e)\approx Exp\left[\Delta A(v_{p})\cdot\frac{\partial\mathcal{M}(Y_{sub}|P_{sub})}{\partial A(v_{p})}\right].$ (6)
 Circuit construction. Edges with scores $|S(e)|$ below a threshold $\tau$ are pruned. The remaining subgraph forms the knowledge circuit $C$. See more details about construction in Appendix B.1
+Circuit in Knowledge Overshadowing.
+In the context of knowledge overshadowing, the metric $\mathcal{M}$ is defined as the difference between the output logits of $Y_{sub}$ and $Y_{dom}$.
+A positive value for $\mathcal{M}$ indicates that the model correctly predicts $Y_{sub}$.
+Therefore, we can represent the circuit where overshadowing occurs as $C_{overshadowed}=\{V,E\}$ with $\mathcal{M}<0$ and $S(E|P_{sub},Y_{dom})\geq\tau$.
+Furthermore, the competition between the dominant and subordinate knowledge pairs during circuit construction is quantified by the activation values of different input $P_{dom}$ and $P_{sub}$, $\Delta A(v_{p})=A_{P_{sub}}(v_{p})-A_{P_{dom}}(v_{p})$.
 3.2.3 Circuit-based Analysis
 PhantomCircuit mainly focuses on the attention heads in $C$ and follows these steps:
-Node Attention Analysis. We identify high attention heads within $C$ by examining their attention scores and patterns, specifically their focus on $\{X_{dom},X_{sub}\}$.
+Node Attention Analysis. We identify high-attention heads within $C$ by examining their attention scores and patterns, specifically their focus on $\{X_{dom},X_{sub}\}$.
 Circuit Structure Analysis.
-We then trace the information flow of these high attention heads by identifying their parent and child nodes to understand their structural role.
+We then trace the information flow of these high-attention heads by identifying their parent and child nodes to understand their structural role.
 Nodes consistently retained in circuits built with different thresholds $\tau$ are also analyzed as key components.
@@
 $\tau_{opt}=\operatorname*{arg\,max}_{\tau}\mathcal{M}(C_{opt}(\tau),P_{sub},Y_{sub}),$ (7)
-where $\mathcal{M}$ measures the circuit’s ability to distinguish $\{X_{dom},X_{sub}\}$, This results in a $C_{opt}(\tau_{opt})$ which is expected to mitigate or eliminate the overshadowing effect for specific input prompts.
+where $\mathcal{M}$ measures the circuit’s ability to distinguish $\{X_{dom},X_{sub}\}$.
+We use a two-stage optimization method to find the optimal number of edges.
+First, we build the circuit by adding edges at regular intervals, revealing that the circuit’s performance exhibits a multi-modal relationship with the edge count.
+Subsequently, we apply the golden-section search algorithm within the optimal range to determine the $\tau_{opt}$ for the $C_{opt}$.
+This results in a $C_{opt}(\tau_{opt})$ which is expected to mitigate or eliminate the overshadowing effect for specific input prompts.
 To automate the overshadowing recovery process and extend its applicability, we simplify the relative pointwise mutual information (R-PMI) method from Zhang et al.
@@
 The $X^{\prime}_{sub}$ yielding the minimum $S_{R\textnormal{-}PMI}$ is identified as $X_{sub}$.
-Furthermore, the $Y_{sub}$ is determined as the token $y_{i}$ from $V_{top}(P_{sub})$ whose average rank improves most significantly when non-subordinate components $X^{\prime}_{sub}$, often the $X_{bg}$, is masked.
+Furthermore, the $Y_{sub}$ is determined as the token $y_{i}$ from $V_{top}(P_{sub})$ whose average rank improves most significantly when non-subordinate components $X^{\prime}_{sub}$, often the $X_{bg}$, are masked.
 $Y_{dom}$ is identified as $y_{i}$ that has the highest average rank across all $P^{\prime}_{sub}$.
@@
 (c) M = 1.4B
-Figure 3: The co-evolution between $\mathcal{RO}$ $\mathcal{LP}$ in different M.
+Figure 4: The co-evolution between $\mathcal{RO}$ & $\mathcal{LP}$ in different M.
 Early high overall loss and low $\mathcal{LP}$ leads to intensive optimization of $P_{dom}$ and $\mathcal{RO}$ rises up to near 100%.
@@
 See more details of our dataset in Appendix C.
-Finetuning Dataset.
-We construct a finetuning dataset to evaluate circuit-base overshadowing recovery method.
+fine-tuning Dataset.
+We construct a fine-tuning dataset to evaluate circuit-base overshadowing recovery method.
 Utilizing virtual knowledge, we preserve the natural language semantics while avoiding prior knowledge editing that could interfere with the natural occurrence of overshadowing.
 For this dataset, we set P = 5 and D = 1M.
-The Appendix D shows the dynamics analysis based on finetuning dataset.
+The Appendix D shows the dynamics analysis based on fine-tuning dataset.
 4.1.2 Models Evaluated
@@
 A uniform learning rate of $10^{-5}$ and batch size of 16 are used for both dataset. Training is conducted on NVIDIA A800 GPUs.
+Figure 5: The circuit-based analysis of proposed PhantomCircuit.
+(a) shows the average attention scores allocated to $\{X_{dom},X_{sub}\}$ across epoch.
+(b) focuses on the onset and recovery phases and shows that the appear and disappear of high-attention heads (attention score greater than 0.2) attribute to the reduce and raise of $\mathcal{RO}$.
+(C) shows the logits and ranks of $Y_{sub},Y_{dom}$ across layers.
+The main structure of circuit (with 400 edges totally) and attention mode show that high-attention head a5.h1 and MLP layer m5 contribute to the juncture of $Y_{sub}$ ’s rank.
+Solid lines denote direct information flow, while dashed lines indicate indirect flow in circuit structure map.
 4.1.4 Evaluation
@@
 The $\mathcal{LP}$ is recorded within each epoch.
-The results are shown in Figures 2 and 3.
-Figure 4: The circuit-based analysis of proposed PhantomCircuit.
-(a) shows the average attention scores allocated to $\{X_{dom},X_{sub}\}$ across epoch.
-(b) focuses on the onset and recovery phases and shows that the appear and disappear of high attention heads (attention score greater than 0.2) attribute to the reduce and raise of $\mathcal{RO}$.
-(C) shows the logits and ranks of $Y_{sub},Y_{dom}$ across layers.
-The main structure of circuit (with 400 edges totally) and attention mode show that high attention head a5.h1 and MLP layer m5 contribute to the juncture of $Y_{sub}$ ’s rank.
-Solid lines denote direct information flow, while dashed lines indicate indirect flow in circuit structure map.
+The results are shown in Figures 3 and 4.
 4.2 Main Result
@@
 Distinctly, a larger D contributes to the earlier onset but also a slower recovery from overshadowing.
-As shown in Figures 2(a) and 2(d) (M=70M, D=2.6M), increasing P (from 5 to 100) significantly shortens or even eliminates the onset phase.
+As shown in Figures 3(a) and 3(d) (M=70M, D=2.6M), increasing P (from 5 to 100) significantly shortens or even eliminates the onset phase.
 This is attributed to more prominent dominant patterns $X_{dom}\leftrightarrow X_{bg}$ being learned and generalized rapidly, even within the first epoch.
 The duration phase also decreases with higher P, because a larger P and a fixed D implies fewer groups of knowledge pairs, leading to less diversity of $P_{sub}$, allowing the model to learn all overshadowed $P_{sub}$ and recover from overshadowing more quickly.
-Figures 2(b) and 2(e) (P=5, D=2.6M) demonstrate that larger models (M) exhibit shorter or absent onset phases, indicating an earlier occurrence of knowledge overshadowing.
+Figures 3(b) and 3(e) (P=5, D=2.6M) demonstrate that larger models (M) exhibit shorter or absent onset phases, indicating an earlier occurrence of knowledge overshadowing.
 This is due to the stronger generalization capabilities of larger models, leading them to quickly learn and overgeneralize dominant patterns.
 However, larger models also show a shortened duration phase and a rapid decline in $\mathcal{RO}$ during recovery, suggesting enhanced capacity to differentiate $\{X_{dom},X_{sub}\}$, thus recovering faster despite earlier overshadowing.
-In Figures 2(c) and 2(f) (P=5, M=70M) larger D leads to an earlier onset of overshadowing, with $\mathcal{RO}$ peaking sooner.
+Figure 6: Overshadowing recovery via optimized circuit.
+In the fine-tuned model, $X_{dom}\leftrightarrow X_{bg}$ causes the original full model to incorrectly predict $Y_{dom}$.
+First, we detect the $X_{sub},Y_{sub},Y_{dom}$ automatically by calculating the minimum $S_{R\textnormal{-}PMI}$ .
+Then, optimizing the knowledge circuit by pruning edges to keep only key nodes enhances the attention on $\{X_{dom},X_{sub}\}$, enabling the recovery from overshadowing and facilitating correct $Y_{sub}$ prediction.
+In Figures 3(c) and 3(f) (P=5, M=70M) larger D leads to an earlier onset of overshadowing, with $\mathcal{RO}$ peaking sooner.
 This is because more data per epoch provides more iterations and exposure to the dominant pattern, accelerating its generalization.
@@
 The increased diversity of $P_{sub}$ in larger datasets requires more epochs for the model to learn all instances and recover.
-Figure 5: Overshadowing recovery via optimized circuit.
-In the finetuned model, $X_{dom}\leftrightarrow X_{bg}$ causes the original full model to incorrectly predict $Y_{dom}$.
-First, we detect the $X_{sub},Y_{sub},Y_{dom}$ automatically by calculating the minimum $S_{R\textnormal{-}PMI}$ Then, optimizing the knowledge circuit by pruning edges to keep only key nodes enhances the attention on $\{X_{dom},X_{sub}\}$, enabling the recovery from overshadowing and facilitating correct $Y_{sub}$ prediction.
 Notably, across various parameter combinations, $\mathcal{RO}$ often approaches 100% in the early training stages and finally recovers to 0%.
@@
 Therefore, the insufficient optimization results in the overshadowing, which is consistent across varying M.
-The occurrence of overshadowing in pretrained LLMs can be understood with dynamics analysis results above.
+The occurrence of overshadowing in pre-trained LLMs can be understood with dynamics analysis results above.
 Initially, large M and D promote the rapid generalization of $X_{dom}\leftrightarrow X_{bg}$ and overshadowing onset.
 The subsequent inadequate optimization of the $P_{sub}$ is exacerbated by the sheer scale and diversity of training data, which brings prolonged overshadowing effect and exceeded sharp recovery effect from large M (e.g., trillion tokens for training Llama-2-7B).
-The knowledge circuit’s attentional allocation to differences between dominant and overshadowed knowledge inputs dictates the extent of knowledge overshadowing.
+The knowledge circuit’s attentional allocation to differences between dominant and overshadowed knowledge inputs influences the extent of knowledge overshadowing.
 Figure 5 (a) shows the variation of circuit’s average attention scores on $\{X_{dom},X_{sub}\}$ throughout the training phase.
@@
 Figure 5 (b) shows the circuit dynamics across the onset and recovery phases of knowledge overshadowing.
-These bar plots show the attention scores of individual attention heads in a specific epoch, indicating that when the $\mathcal{RO}$ declines, some attention heads, defined as high attention heads, exhibiting high focus on $\{X_{dom},X_{sub}\}$ emerge within the circuit.
-The threshold for high attention is set at 0.2 for the length of $X_{sub}$ is one fifth of the length of input prompt in synthetic dataset.
+These bar plots show the attention scores of individual attention heads in a specific epoch, indicating that when the $\mathcal{RO}$ declines, some attention heads, defined as high-attention heads, exhibiting high focus on $\{X_{dom},X_{sub}\}$ emerge within the circuit.
+The threshold for high-attention is set at 0.2 for the length of $X_{sub}$ is one fifth of the length of input prompt in synthetic dataset.
 Conversely, when knowledge overshadowing intensifies, a subset of these critical attention heads tends to disappear from the circuit.
+Table 1: Ablation Study for high-attention Heads on fine-tuning Dataset. The model’s performance degrades with the ablation of high-attention heads, which validates the circuit’s faithfulness.
+Model Proportion of Ablated High Attn. Heads 10% 20% 50%
+gpt2-medium Performance Degradation 1.03 1.43 1.67
+Attn. Score on $X_{sub}$ Degradation 0.044 0.086 0.141
+pythia-410m Performance Degradation 0.92 1.30 1.42
+Attn. Score on $X_{sub}$ Degradation 0.039 0.074 0.126
 Furthermore, by focusing on the circuit’s internal mechanisms and structure within a specific epoch, as shown in Figure 5 (c), we leverage the circuit-based analysis of our proposed PhantomCircuit.
@@
 Subsequently, by focusing on the circuit structure, we identify the internal mechanisms driving this juncture.
-A high attention head, a5.h1 (layer 5), crucially channels information to the subsequent MLP layer, m5.
-The a5.h1 also appears to inherit its high attention state from earlier layers, mediated by the layer 4 MLP (m4), leading it to not only continue elevating the logits for $Y_{sub}$ but also to attenuate the previously rapid growth of $Y_{dom}$’s logits, thereby facilitating the observed rank reversal.
+A high-attention head, a5.h1 (layer 5), crucially channels information to the subsequent MLP layer, m5.
+The a5.h1 also appears to inherit its high-attention state from earlier layers, mediated by the layer 4 MLP (m4), leading it to not only continue elevating the logits for $Y_{sub}$ but also to attenuate the previously rapid growth of $Y_{dom}$’s logits, thereby facilitating the observed rank reversal.
 The attention map on the right confirms its high focus on $\{X_{dom},X_{sub}\}$.
@@
 Figure 6 illustrates our findings.
-We first finetune the model on finetuning dataset.
+We first fine-tune the model on fine-tuning dataset.
 During the recovery phase, we randomly choose some $P_{sub}$.
@@
 With these components, we obtain the $P_{dom}$ as corrupt input and $P_{sub}$ as clean input to construct a knowledge circuit.
-A golden section search algorithm is employed to determine the optimal number of edges for building $C_{opt}$.
-The optimized circuit structure map shows that some attention heads are pruned, which are often low attention heads, or exhibit no significant attentional pattern towards the core task-relevant information.
-Retained high attention heads are key to differentiating $P_{dom}$ from $P_{sub}$ which give significant attention to $\{X_{dom},X_{sub}\}$.
+The optimized circuit graph shows that some attention heads are pruned, which are often low attention heads, or exhibit no significant attentional pattern towards the core task-relevant information.
+Retained high-attention heads are key to differentiating $P_{dom}$ from $P_{sub}$ which give significant attention to $\{X_{dom},X_{sub}\}$.
 Some low attention heads also remain, implying that even in the circuit, processing background knowledge $X_{bg}$ and linking it to the distinctive elements $X_{sub}$ is crucial for correct inference.
@@
 Finally, the circuit successfully produces $Y_{sub}$, demonstrating the elimination of the overshadowing effect.
-Future work will enhance the circuit optimization metric $\mathcal{M}$ by incorporating $Y_{sub}$’s absolute logit alongside the logit difference with $Y_{dom}$ for more effective guidance.
-Developing a comprehensive evaluation framework for circuit-based recovery is also crucial.
-These steps will evolve PhantomCircuit into an integrated platform for efficient analysis and robust optimization of knowledge overshadowing.
+4.3 Circuit Faithfulness Study
+Based on the work of Marks et al.
+(2024); Hanna et al.
+(2024), the EAP-IG method builds circuits with high faithfulness and efficiency, ensuring our findings based on circuit analysis are reliable.
+To verify this key property in the context of knowledge overshadowing, we conduct an ablation study on high-attention heads.
+We first construct optimal circuits for pythia-410m and gpt2-medium models trained on the fine-tuning dataset.
+Subsequently, we removed the top 10%, 20%, and 50% of attention heads by score and observed the impact on performance metric $\mathcal{M}$ and attention scores, with results shown in Table 1.
+The results show that ablating just 10% of the high-attention heads leads to a significant drop in both the attention scores on $X_{dom},X_{sub}$} and the circuit performance metric $\mathcal{M}$.
+This performance degradation is substantial enough to alter the output in many cases, preventing the model from accurately predicting $Y_{sub}$.
+In contrast, at 50% ablation, the drop in $\mathcal{M}$ becomes less pronounced because the model’s ability to predict both the $Y_{sub}$ and $Y_{dom}$ severely declines.
+These findings validate the critical impact of the nodes in optimal circuits, especially high-attention heads, on model performance and confirm the circuit’s faithfulness.
 5 Conclusion
@@
 PhantomCircuit first analyzes the training dynamics of overshadowing, finding that dominant knowledge popularity, model size, and dataset size critically shape the onset, duration, and recovery of overshadowing.
-Apart from that, the persistent overshadowing in pretrained models stems from inadequately optimized subordinate knowledge loss.
+Apart from that, the persistent overshadowing in pre-trained models stems from inadequately optimized subordinate knowledge loss.
 By analyzing knowledge circuits, we find that changes in critical attention heads’ focus on subordinate knowledge directly correlate with the recovery or onset of overshadowing.
@@
   1. 1.
-    The dynamic analysis of knowledge circuits throughout training is computationally intensive, potentially hindering scalability to very large models or extensive training.
-We aim to develop more computationally efficient techniques for approximating circuit evolution, such as checkpoint-based analysis or lightweight probing.
+    Despite of the experiments on gpt2-medium and natural data case (Appendix E), this study’s findings primarily rely on the Pythia model suite and synthetic/fine-tuning datasets, which are chosen to ensure architectural consistency and data purity for controlled analysis.
+It remains to be validated whether these conclusions can be directly generalized to other model architectures and more complex, noisy real-world data.
   2. 2.
-    This study concentrates on a specific type of knowledge overshadowing, leaving more complex or subtle interference patterns unaddressed.
-Future work will broaden PhantomCircuit’s scope to investigate a wider range of overshadowing phenomena, including those in multi-hop reasoning.
+    The current work concentrates on direct knowledge interference arising from imbalanced knowledge popularity.
+The applicability of PhantomCircuit to practical tasks like multi-hop reasoning and code generation, remains a key area for future exploration.
   3. 3.
-    Future efforts will focus on evolving our instance-specific circuit optimizations into a generalized mitigation toolkit, supported by a comprehensive evaluation framework.
-Key improvements will target the precision of automated overshadowed knowledge identification and the broader efficacy of circuit-based interventions.
-Ultimately, we aim to develop PhantomCircuit as a robust platform for both in-depth analysis and effective, generalizable overshadowing mitigation.
+    Performing dynamic analysis of knowledge circuits throughout the entire training process is computationally intensive, which could potentially hinder its scalability to very large models or extensive training runs.
+Future work aims to develop more computationally efficient techniques, such as more efficient circuit extraction methods like Information Flow Routes (IFR), to mitigate this challenge.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (Grant No.62506318); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); CAAI-Ant Group Research Fund; Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
 Appendix A More Related Work
@@
 (2023).
-Pretrained on large-scale text corpora, LLMs have acquired extensive real-world knowledge from web sources.
+pre-trained on large-scale text corpora, LLMs have acquired extensive real-world knowledge from web sources.
 As a result, models such as InternLM2.5 Cai et al.
@@
 The process begins by representing the LLM as a directed acyclic graph (DAG), $G=(V,E)$, where $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, and $E$ represents the information flow between these components.
-The goal is to identify a subgraph $C\subseteq G$ that is critical for recognize the key component of a given input prompt, particularly in knowledge overshadowing, is $X_{dom}$ and $X_{sub}$, the difference between $P_{dom}$ and $P_{sub}$.
-The adapted construction method is similar to edge attribution patching(EAP) Conmy et al. (2023), which involves:
+Our goal is to identify a subgraph $C$ that is critical for recognizing the key component of a given input prompt.
+In the context of knowledge overshadowing, this component is $\{X_{dom},X_{sub}\}$, which represents the difference between $P_{dom}$ and $P_{sub}$.
+The adapted construction method is similar to Edge Attribution Patching with Integrated Gradients (EAP-IG) Hanna et al. (2024), which involves:
   1. 1.
@@
     where $Y_{target}$ is ideally $Y_{sub}$.
-The expectation $\mathbb{E}$ is taken over instances of $P_{sub}$ in our evaluation set $Z_{sub}$.
+The expectation $\mathbb{E}$ is taken over instances of $P_{sub}$ in evaluation set.
 In practice, methods like Integrated Gradients (IG) are often used to refine this attribution by integrating gradients along a path from a baseline (corrupted) input to the actual (clean) input.
@@
     This constructed circuit $C_{sub}$ is then analyzed to understand how dominant knowledge $K_{dom}$ might overshadow $K_{sub}$ by examining the attentional features and information flow within it, especially when processing $P_{sub}$.
+  5. 5.
+    Threshold $\tau$ Optimization: By constructing a circuit for a given edge pruning threshold, $\tau$, and then evaluating its performance metric, $M$, we can establish the relationship between $\tau$ and $\mathcal{M}$.
+Therefore, optimizing the circuit is equivalent to finding an optimal threshold, $\tau_{opt}$, that maximizes the performance metric $\mathcal{M}$.
+In practice, however, we directly optimize the number of edges $n$ rather than the threshold $\tau$.
+We divide this optimization process into two stages.
+In the first stage, we start from an initial edge count and incrementally increase the number of edges at uniform intervals to map out the relationship between $\mathcal{M}$ and $n$.
+We find that this relationship is multi-modal, as shown in Figure 8.
+In the second stage, we employ the golden-section search algorithm within the most promising interval identified previously to precisely locate the optimal number of edges, $n_{opt}$.
+This process yields the optimized circuit, $C_{opt}(n_{opt})$.
 B.2 Automated Component Identification for Recovery
@@
 Summing only the negative R-PMI values to obtain
-$S_{R\textnormal{-}PMI^{-}}(P_{sub},P^{\prime}_{sub})=\sum\min(R\textnormal{-}PMI(y_{i}),0).$
-The $X^{\prime}_{sub}$ that yields the minimum (most negative) $S_{R\textnormal{-}PMI^{-}}$ is identified as the primary overshadowed component, $X_{sub}^{*}$.
+$S_{R\textnormal{-}PMI}(P_{sub},P^{\prime}_{sub})=\sum\min(R\textnormal{-}PMI(y_{i}),0).$
+The $X^{\prime}_{sub}$ that yields the minimum (most negative) $S_{R\textnormal{-}PMI}$ is identified as the primary overshadowed component, $X_{sub}^{*}$.
 This selection is based on the rationale that removing the true $X_{sub}^{*}$ most strongly exposes the model’s bias towards outputs favored by the dominant knowledge pattern.
+In practical implementation, we enhance the identification of $X_{sub}^{*}$ by calculating a weighted $S_{R\textnormal{-}PMI}$.
+To do this, we weight each token’s $R-PMI(y_{i})$ value by the variance of its log probability change, $var(y_{i})$, observed when masking different candidates of $X_{sub}$.
+This results in $S_{R\textnormal{-}PMI}(P_{sub},P^{\prime}_{sub})_{weighted}=\sum\min(R\textnormal{-}PMI(y_{i})\cdot Var(y_{i}),0)$, which amplifies the effect from the $Y_{s}ub$ token which is most sensitive to the masking of $X^{*}_{sub}$, enabling a more robust identification.
 Identifying Target Subordinate Output $Y_{sub}$.
 The intended subordinate output $Y_{sub}$ is identified by assessing which token from $V_{top}(P_{sub})$ (the top-k candidates for the original prompt $P_{sub}=(X_{bg},X_{sub})$) exhibits the most significant improvement when the overshadowing influence of background knowledge ($X_{bg}$) or other non-subordinate components is mitigated.
-Specifically, we generate contrastive prompts $P^{\prime}_{sub}$ by masking or altering components of $X_{bg}$ (or other identified non-subordinate elements that contribute to the $X_{bg}\leftrightarrow Y_{dom}$ association) within $P_{sub}$.
-$Y_{sub}$ is then the token $y_{i}\in V_{top}(P_{sub})$ that shows the most substantial rank improvement (or largest increase in log probability) in these modified prompts $P^{\prime}_{sub}$ compared to its rank in the original $P_{sub}$.
-This rank elevation signifies the “unmasking” of the true subordinate answer as the dominant, overshadowing associations are weakened.
+Specifically, we generate contrastive prompts $P^{\prime}_{sub}$ by masking or altering components of $X_{bg}$ within $P_{sub}$.
+$Y_{sub}$ is then the token $y_{i}\in V_{top}(P_{sub})$ that shows the most substantial rank improvement in these modified prompts $P^{\prime}_{sub}$ compared to its rank in the original $P_{sub}$.
+Furthermore, to properly value the rank improvements of initially high-ranked tokens like $Y_{dom}$ and $Y_{sub}$, which have limited room for elevation, we introduce a weight to measure the original rank’s impact.
+This yields the final weighted rank elevation metric for every predicted top-$k$ next-tokens: $\text{Weighted Rank Elevation}=(k-\text{rank}_{\text{original}})\times\text{Avg.
+Elevation of Rank}.$
+This approach preserves the effect of rank improvements for top-ranked tokens.
+And the rank elevation signifies that the mitigation of impact from $X_{dom}\leftrightarrow X_{bg}$ make overshadowing association weaken, which improves the rank of $Y_{sub}$.
 Identifying Dominant Output $Y_{dom}$.
@@
 This token represents the model’s most consistent, default output tendency when specific subordinate cues are variably weakened, likely reflecting the pervasive influence of dominant knowledge associated with the background $X_{bg}$.
-[…]
+With $X_{bg}$ (background knowledge), $X_{sub}$ (identified subordinate component), the expected $Y_{sub}$, and the interfering $Y_{dom}$ established, we prepare the paired inputs required for knowledge circuit construction.
+The clean input is the original subordinate prompt $P_{sub}=(X_{bg},X_{sub})$, for which the desired output is $Y_{sub}$.
+To create the corrupt input $P_{dom}$, which is designed to elicit the overshadowing effect and output $Y_{dom}$, we maintain the background knowledge $X_{bg}$ but replace the subordinate component $X_{sub}$ with a generic placeholder token, such as ’something’.
+Thus, $P_{dom}=(X_{bg},\text{``something''})$.
+This specific formulation of $P_{dom}$ ensures that while the input structure is similar to $P_{sub}$, the absence of $X_{sub}$ allows the strong $X_{bg}\leftrightarrow Y_{dom}$ association to dominate, leading to the incorrect prediction $Y_{dom}$.
+These paired inputs, $P_{sub}$ and $P_{dom}$, then serve as the foundation for the activation difference calculations in our circuit analysis.
 Some more circuit-based overshadowing recovery cases are shown in Table 2.
 Table 2: Circuit-based overshadowing recovery cases
-{tblr}
-[…]
-[…]
+Case $P_{sub}$ with $\{X_{sub}\}$ $\mathcal{M}$ indicator (logits difference) $Y_{dom}\;\&\;Y_{sub}$ Full Model Top 5 Prediction Circuit Top 5 Prediction
+Case 1 Analysis of the Chrono-Filter device efficiency for temporal sorting shows outcome {filtration overload} Original model:-1.283 & Circuit: 0.764 |Time| & |Tem| Rank 0: Logit: 16.18 Prob: 32.18% Token: |Tem| Rank 0: Logit: 21.03 Prob: 73.70% Token: |Time|
 Rank 1: Logit: 15.42 Prob: 14.98% Token: |Time| Rank 1: Logit: 19.75 Prob: 20.42% Token: |Tem|
@@
 Rank 4: Logit: 13.84 Prob: 3.10% Token: |Custom| Rank 4: Logit: 16.29 Prob: 0.64% Token: |Sp|
-[…]
-[…]
+Case 2 Constructing psionic wave emitters necessitates precise tuning involving specialized harmonic {feedback loop} Original model: -0.745& Circuit: 2.490 |Wave| & |E| Rank 0: Logit: 17.64 Prob: 37.18% Token: |Wave| Rank 0: Logit: 17.96 Prob: 46.34% Token: |E|
 Rank 1: Logit: 16.89 Prob: 17.65% Token: |E| Rank 1: Logit: 16.47 Prob: 10.39% Token: |Ps|
@@
 Rank 4: Logit: 15.15 Prob: 3.09% Token: |emitter| Rank 4: Logit: 14.89 Prob: 2.15% Token: |emitter|
-[…]
-[…]
+Case 3 Analyzing Ectoplasmic Conduit energy transfer efficiency through degrading {structure reinforcement reveals} Original model: -3.019 & Circuit: 0.523 |Transfer| & |Emit| Rank 0: Logit: 49.08 Prob: 49.26% Token: |Transfer| Rank 0: Logit: 44.54 Prob: 47.18% Token: |Emit|
 Rank 1: Logit: 48.68 Prob: 33.09% Token: |St| Rank 1: Logit: 44.01 Prob: 27.96% Token: |Transfer|
@@
 Rank 4: Logit: 46.06 Prob: 2.41% Token: |Emit| Rank 4: Logit: 41.64 Prob: 2.61% Token: |Mi|
-[…]
-[…]
+Case 4 Shard Relic residual energy output response to sudden energy {conduit field spikes} shows Original model: -2.623 & Circuit: 3.202 |Output| & |St| Rank 0: Logit: 40.96 Prob: 91.94% Token: |Output| Rank 0: Logit: 34.61 Prob: 51.02% Token: |St|
 Rank 1: Logit: 38.34 Prob: 6.67% Token: |St| Rank 1: Logit: 34.24 Prob: 35.19% Token: |Field|
@@
 Appendix C Dataset Details
+(a) The $\mathcal{RO}$ during training phase of the different fine-tuning dataset size (D).
+(b) The co-evolution of $\mathcal{LP}$ and $\mathcal{RO}$ during training phase on fine-tuning dataset.
+Figure 7: The dynamics analysis of knowledge overshadowing in fine-tuning dataset.
 C.1 Detailed Synthetic Dataset Construction
@@
 Table 3: Illustrative examples from the synthetic dataset (P=5). Each data entry is a row, with fine lines separating entries within a group. Token IDs are shown.
-{tblr}
-[…]
-[…]
+Group $X_{bg}$ $X_{dom}$ $Y_{dom}$ $X_{sub}$ $Y_{sub}$
+Group 1 $[10030,16936,1050,10565]$ $10279$ $20730$
 $[10030,16936,1050,10565]$ $24327$ $20730$
@@
 $[10030,16936,1050,10565]$ $18941$ $3519$
-[…]
+Group 2 $[17026,8837,3802,28741]$ $2496$ $1077$
 $[17026,8837,3802,28741]$ $3530$ $1077$
@@
 $[17026,8837,3802,28741]$ $25814$ $5374$
-[…]
+Group 3 $[18131,14501,21161,311]$ $4706$ $7790$
 $[18131,14501,21161,311]$ $778$ $7790$
@@
 $[18131,14501,21161,311]$ $8447$ $6129$
-C.2 Finetuning dataset
-For the finetuning dataset, we utilized the Qwen-Long API to generate instances of virtual knowledge.
+C.2 fine-tuning dataset
+For the fine-tuning dataset, we utilized the Qwen-Long API to generate instances of virtual knowledge.
 This generated data subsequently underwent manual review to identify and remove any instances that are overly repetitive or semantically too similar, ensuring a degree of diversity, resulted in D = 1M.
 A key distinction from the synthetic dataset construction is that we did not strictly control token lengths for each component in this dataset.
-Instead of randomly sampled token IDs, the finetuning dataset consists of actual linguistic statements that, while syntactically and semantically coherent, represent virtual (i.e., fabricated but plausible) knowledge.
+Instead of randomly sampled token IDs, the fine-tuning dataset consists of actual linguistic statements that, while syntactically and semantically coherent, represent virtual (i.e., fabricated but plausible) knowledge.
 The underlying pattern of dominant and subordinate knowledge construction, however, mirrors that of the synthetic dataset.
 As an example of this dataset, we set the knowledge popularity P=5. Some illustrative cases from the dataset are shown in Table 4.
-Table 3: Illustrative examples from the Finetuning Dataset (P=5). Each data entry is a row, with fine lines separating entries within a group.
-{tblr}
-[…]
-[…]
+Table 4: Illustrative examples from the fine-tuning Dataset (P=5). Each data entry is a row, with fine lines separating entries within a group.
+Group $X_{bg}$ $X_{dom}$ $Y_{dom}$ $X_{sub}$ $Y_{sub}$
+Group 1 Analysis of the Starfire Crystal Engine reveals primary energy output peak resonance Pure Nova
 Analysis of the Starfire Crystal Engine reveals primary energy output idle cycle Pure Nova
@@
 Analysis of the Starfire Crystal Engine reveals primary energy output cold ignition Quantum Flux
-[…]
+Group 2 Observation of the Chronos Orb’s temporal distortion signature indicates full moonlight Stable Chrono-displacement
 Observation of the Chronos Orb’s temporal distortion signature indicates near Leyline Stable Chrono-displacement
@@
 Observation of the Chronos Orb’s temporal distortion signature indicates sudden impact Temporal Anomaly
-[…]
+Group 3 Study of the Xylosian biology reveals primary nutrient absorption method consuming light Molecular Assimilation
 Study of the Xylosian biology reveals primary nutrient absorption method after sonic bath Molecular Assimilation
@@
 Study of the Xylosian biology reveals primary nutrient absorption method xenoflora consumption Crystalline Excretion
-(a) The $\mathcal{RO}$ during training phase of the different finetuning dataset size (D).
-(b) The co-evolution of $\mathcal{LP}$ and $\mathcal{RO}$ during training phase on finetuning dataset.
-Figure 6: The dynamics analysis of knowledge overshadowing in finetuning dataset.
-Appendix D More Dynamics Analysis on Finetuning Dataset
-In addition to validating the efficacy of our circuit-based overshadowing recovery method, the finetuning dataset serves a dual purpose.
+(a) gpt2-medium
+(b) pythia-410m
+Figure 8: The relationship between circuit performance, high-attention heads proportion, average attention score on $X_{sub}$ and the number of edges in gpt2-medium and pythia-410m.
+Appendix D More Dynamics Analysis on fine-tuning Dataset
+In addition to validating the efficacy of our circuit-based overshadowing recovery method, the fine-tuning dataset serves a dual purpose.
 We also leverage it to empirically verify our conclusions regarding the training dynamics of knowledge overshadowing, specifically concerning the impact of Dataset Size (D).
 Consistent with our dynamic analysis findings, we investigate whether a larger D indeed correlates with a slower recovery rate from knowledge overshadowing and a prolonged duration of the hallucination effect.
-To this end, we conduct experiments on the finetuning dataset by fixing Knowledge Popularity at P=5 and Model Size at M=70M, while varying D across values of 0.1M, 0.5M, and 1M tokens.
+To this end, we conduct experiments on the fine-tuning dataset by fixing Knowledge Popularity at P=5 and Model Size at M=70M, while varying D across values of 0.1M, 0.5M, and 1M tokens.
 The results, as depicted in Figure 7(a), corroborate this relationship.
-Furthermore, under the specific configuration of P=5, M=70M, and D=0.5M on the finetuning dataset, we re-examine the interplay between the loss proportion of subordinate knowledge ($\mathcal{LP}$) and the relative overshadowing rate ($\mathcal{RO}$).
+Furthermore, under the specific configuration of P=5, M=70M, and D=0.5M on the fine-tuning dataset, we re-examine the interplay between the loss proportion of subordinate knowledge ($\mathcal{LP}$) and the relative overshadowing rate ($\mathcal{RO}$).
 As shown in Figure 7(b), the observations again support the hypothesis that insufficient optimization of subordinate knowledge contributes to the persistence of knowledge overshadowing.
-It is noteworthy that distinct behaviors are observed when comparing the finetuning dataset to the synthetic dataset.
-Firstly, the recovery from overshadowing on the finetuning dataset is generally slower than on the synthetic dataset for same D.
-This can be attributed to the richer semantic relationships and greater complexity inherent in the natural language of the finetuning data, which presents a more challenging learning task.
-Secondly, we observe that the finetuning dataset exhibits a minimal or absent onset phase for knowledge overshadowing, where $\mathcal{RO}$ typically rise.
-This is because finetuning commences from a pretrained model, which has already moved beyond the initial epochs of chaotic, random predictions.
-Consequently, the model can very rapidly generalize strong association patterns present in the finetuning data.
-Moreover, the diverse and varied forms of data within the finetuning set may act akin to a beneficial noise signal, prompting the model to pay closer attention to distinguishing features and differences.
+It is noteworthy that distinct behaviors are observed when comparing the fine-tuning dataset to the synthetic dataset.
+Firstly, the recovery from overshadowing on the fine-tuning dataset is generally slower than on the synthetic dataset for same D.
+This can be attributed to the richer semantic relationships and greater complexity inherent in the natural language of the fine-tuning data, which presents a more challenging learning task.
+Secondly, we observe that the fine-tuning dataset exhibits a minimal or absent onset phase for knowledge overshadowing, where $\mathcal{RO}$ typically rise.
+This is because fine-tuning commences from a pre-trained model, which has already moved beyond the initial epochs of chaotic, random predictions.
+Consequently, the model can very rapidly generalize strong association patterns present in the fine-tuning data.
+Moreover, the diverse and varied forms of data within the fine-tuning set may act akin to a beneficial noise signal, prompting the model to pay closer attention to distinguishing features and differences.
 This inherent data diversity can help preemptively mitigate or even eliminate the early onset stage of knowledge overshadowing that might otherwise be observed.
+Appendix E More Analysis on Natural Language Data
+To further validate the generalization of our findings, we conduct experiments on the pre-trained Pythia-410m and gpt2-medium models using natural language prompts.
+The paired inputs for this test are: $P_{dom}$: "The name of this North Korean politician is", $Y_{dom}$: "Kim Jong Un", $X_{dom}$: "politician"} and $P_{sub}$: "The name of this North Korean singer is", $Y_{sub}$: "Hyon Song-wol", $X_{sub}$: "singer"}.
+Since it is hard to reproduce the pre-training process of these models, we instead used the number of edges in the circuit as the independent variable.
+We investigated the relationship among three metrics: circuit performance, the proportion of high-attention heads, and the average attention score on $X_{sub}$.
+The results are shown in Figure 8, which lead to the following conclusions:
+  • The model’s overall attention to the different parts $\{X_{dom},X_{sub}\}$ increases as the number of edges is reduced through optimization.
+Correspondingly, the model’s performance improves with this rise in attention.
+This validates our conclusion that higher attention on $\{X_{dom},X_{sub}\}$ is key to mitigating knowledge overshadowing.
+  • The proportion of high-attention heads also increases during edge optimization.
+This supports the conclusion that retaining these high-attention heads plays a crucial role in directing the model’s focus toward $\{X_{dom},X_{sub}\}$.
+  • The model’s performance exhibits a notable multi-modal relationship with the number of edges, which makes the optimization of the circuit’s edge count a more complex task.
+Additionally, our preliminary exploration suggests that knowledge overshadowing can also occur in multi-hop reasoning.
+When the pre-trained OLMo-7B-Instruct model is given the prompt: "The old country of this famous scientist who is not famous for theory of relativity is", it outputs "Germany".
+Furthermore, with the prompt: "The name of this scientist who is not famous for theory of relativity is", the model outputs "Albert".
+We can infer from this that the model still defaults to "Albert Einstein" as the basis for its multi-hop reasoning.